OpenStudy (anonymous):

What is the solution of the system? Use substitution. y = x + 6 y = 2x A. (6, 12) B. (–12, –6) C. (–6, –12) D. (2, 4)

OpenStudy (anonymous):

how do u do that

OpenStudy (dan815):

oh i see i think i read your question wrong it should be y=x+6 y=2x right

OpenStudy (anonymous):

ye

OpenStudy (dan815):

put that x+6 where y is and solve for x

OpenStudy (dan815):

x+6=2x... solve for x

OpenStudy (aric200):

ITS A

OpenStudy (aric200):

X = 6

OpenStudy (aric200):

I know I'm a bum for doing that :O

OpenStudy (anonymous):

tnx aric u saved alot of time

OpenStudy (aric200):

Yeah I understand. Do you know how to do it and we are being lazy though or do you need to know?

OpenStudy (anonymous):

The school cafeteria sells two kinds of wraps: vegetarian and chicken. The vegetarian wrap costs $1.00 and the chicken wrap costs $1.80. Today they made $98.80 from the 70 wraps sold. How many of the wraps sold were vegetarian? A. 36 wraps B. 37 wraps C. 30 wraps D. 34 wraps